# Statistical Taylor Expansion: A New and Path-Independent Method for Uncertainty Analysis 111

### v12

@@ 9.10 Using Library Sine for Linear Signal
As shown by the difference between Figure 34 and 35, Linear signals computed using the Library sine function can introduce unspecified numerical errors up to $10^{3}$ times larger into the results through the Library $\cos(x)/sin(x)$. The key question is whether variance arithmetic can effectively track these additional errors.
@@ 9.11 Ideal Coverage
Adding sufficient noise to the input can suppress unspecified input errors and thereby achieve ideal coverage. For example, applying an input noise of $10^{-3}$ to a Linear signal when using the Library sine function result in ideal coverage.
Figure 50 illustrates the error deviations and uncertainty means:
@@
  • As expected, the result uncertainty means for the reverse transformations decrease with FFT order $L$ as $\sqrt{1/2}^{L}$.
@@
  • As expected, the result error deviations for the forward and reverse transformations remain constant at $1$, whereas those for the roundtrip transformation decay exponentially toward 0 as the FFT order increases.
@@
The range of ideal coverage depends on how accurately the specified input uncertainty represents the actual input noise. For Linear signals computed using the Library sine function, Figure 52 and 53 present the error deviation versus the added noise and FFT order for the forward and reverse transformations, respectively. Ideal coverage corresponds to the region where the error deviation equals $1$. Outside this region, proper coverage cannot be achieved. Because uncertainties grow more slowly in the reverse transformation than in the forward transformation, the reverse transformation exhibits a smaller ideal coverage region. Furthermore, as numerical errors increase with computational load, the range of input noise that produces ideal coverage decreases with increasing FFT order. At sufficiently high FFT orders, visually beyond FFT order $25$ for the reverse transformation, ideal coverage may no longer be achievable. Although FFT is widely regarded as one of the most robust numerical algorithms, and generally insensitive to input errors, it can still fail due to numerical errors in the Library sine function. Such deterioration in calculation accuracy is not easily detectable when using conventional floating-point arithmetic.
In contrast, for Linear signals computed using the Quart sine functions, Figure 54 and 55 show the results for the forward and reverse transformations, respectively. The ideal coverage region is significantly larger, and proper coverage is also achieved in adjacent regions. The forward transformations exhibits a broader ideal coverage region than the reverse transformations.
@@ 9.12 Prec Sine Function
Signal vs Sine Transformation Quart Prec Library
Sin/Cos Forward $10^{-15}$ $10^{-15}$ $10^{-14}$
Sin/Cos Reverse $10^{-12}$ $10^{-12}$ $10^{-11}$
Linear Forward $10^{-9}$ $10^{-9}$ $10^{-6}$
Linear Reverse $10^{-7}$ $10^{-7}$ $10^{-3}$
Table 2: The minimal required noise to achieve ideal coverage for FFT transformations at FFT order $18$ for different sine functions.
The Quart sine function tends to overestimate the uncertainty, while the Library sine function underestimates it. This raises the question of whether a sine function with a more accurate uncertainty can be developed. If the Quart sine function is constructed using the 128-bit floating-point library of gcc and then converted to 64-bit values, the difference between each 128-bit value and its corresponding 64-bit value can be used as the uncertainty for the 64-bit representation. The resulting indexed sine function is referred to as the Prec sine function.
Using $\sin(x)^{2}+\cos(x)^{2}-1$ test to compare the Prec and the Quart sine functions, Figure 56 shows that the the Prec sine function produces slightly larger value errors, while having half uncertainties. Figure 57 and 58 demonstrate that the the Prec sine function slightly underestimates uncertainties, maintaining a stable error deviation around $1.8$ for both forward and reverse transformations without added noise. Compared with Figure 55, Figure 59 shows that the ideal coverage achieved using the the Prec sine function is identical to that obtained with the Quart sine function. The observation that the 128-bit sine function provides no noticeable improvement over the 64-bit sine function raises concerns about the statistical legitimacy of directly using the difference as the uncertainty, the accuracy of the gcc 128-bit numerical library, or the algorithm used to convert a 128-bit floating-point values to 64-bit.
Table 2 compares using different sine functions for Sin/Cos or Linear signals. Both Quart and the the Prec sine function are not ideal. An ideal sine function should have error deviations closer to $1$ without added noise.
@@ 9.13 Summary
The Library sine functions implemented using conventional floating-point arithmetic have been shown to contain numerical errors equivalent up to $10^{-3}$ of the input precision in the worst cases of FFT transformations. These $\sin(x)$ numerical errors increase periodically with $x$, potentially resonating with periodic input signals. Computational result obtained with floating-point arithmetic are highly sensitive to the input data, the scale of computation, and the specific implementation of the mathematical library. This implies that a small-scale test cannot reliably qualify or predict the behavior of large-scale calculations. The impact of numerical errors within mathematical libraries has not received sufficient attention. Moreover, such errors in floating-point arithmetic are inherently difficult to detect and quantify.
@@ 11.3 Improvements Needed
Bound momentum $\zeta(2n)$ needs to be extended to all probability distributions. Its asymptotic behavior $2n\rightarrow+\infty:\zeta(2n)\rightarrow\kappa^{2n}/(2n)$ needs to be generalized.
@@
Determining the uncertainty upper bounds for each analytic expression through analytic methods is an important next step. These results can validate variance arithmetic, and provide performance enhancement to variance arithmetic when the convergence of an analytic expression has to be deduced numerically from the full expansion each time.
@@
As an enhancement to dependency tracing, source tracing identifies each input’s contribution to the overall result uncertainty. This capability enables engineers to pinpoint the primary sources of measurement inaccuracy and in turn guide targeted improvements in data acquisition and processing strategies.
A key open question is whether variance arithmetic can be adapted to achieve ideal coverage for floating-point rounding errors. Once the mechanism of rounding error is thoroughly understood, developing variance arithmetic with ideal coverage for such errors would be quite valuable because many theoretical calculations lack explicit input uncertainties.
@@ 11.4 Acknowledgments
12 Statements and Declarations
Competing Interests: The authors have no competing interests to declare that are relevant to the content of this article.
13 Data Availability Statement
The datasets used in this study are all generated in the open-source project at https://github.com/Chengpu0707/VarianceArithmetic. The execution assistance and explanation of the above code are available from the author upon reasonable request.

### v13

@@ 9.11 Using Library Sine for Linear Signal
As shown by the difference between Figure 34 and 35, Linear signals computed using the Library sine function can introduce unspecified numerical errors up to $10^{3}$ times larger into the results through the Library $\cos(x)/sin(x)$. The key question is whether variance arithmetic can effectively track these additional unspecified errors.
@@ 9.12 Ideal Coverage
Code Signal Sine function Forward Reverse Roundtrip
Python Sin/Cos Quart $10^{-16}$ $10^{-12}$ $10^{-14}$
Java, C++ Sin/Cos Quart $10^{-16}$ $10^{-12}$ $10^{-14}$
Python Sin/Cos Library $10^{-16}$ $10^{-11}$ $10^{-12}$
Java, C++ Sin/Cos Library $10^{-16}$ $10^{-11}$ $10^{-13}$
Python Linear Quart $10^{-11}$ $10^{-7}$ $10^{-8}$
Java, C++ Linear Quart $10^{-11}$ $10^{-7}$ $10^{-8}$
Python Linear Library $10^{-11}$ $10^{-3}$ $10^{-8}$
Java, C++ Linear Library $10^{-11}$ $10^{-3}$ $10^{-8}$
Table 2: The measured minimal required noise to achieve ideal coverage for FFT transformations at FFT order $18$ for different signals and sine functions.
Adding noise to the input can suppress unspecified input errors and improve the portion of specified versus unspecified input uncertainties. Table 2 shows the measured minimal required noise to achieve ideal coverage for FFT transformations at FFT order $18$ for different signals and sine functions.
Figure 54 illustrates the error deviations and uncertainty means when applying an input noise of $10^{-3}$ to a Linear signal when using the Library sine function result in ideal coverage for all the transformations:
@@
  • As expected, the result uncertainty means for the reverse transformations decrease with FFT order $L$ as $1/\sqrt{2}^{L}$.
@@
  • As expected, the result error deviations for the forward and reverse transformations remain constant at $1$.
@@
The range of ideal coverage depends on how accurately the specified input uncertainty represents the actual input noise. For Linear signals computed using the Library sine function, Figure 56 and 57 present the error deviation versus the added noise and FFT order for the forward and reverse transformations, respectively. Ideal coverage corresponds to the region where the error deviation equals $1$. Outside this region, proper coverage cannot be achieved for the reverse transformation. Because uncertainties grow more slowly in the reverse transformation than in the forward transformation, the reverse transformation exhibits a smaller ideal coverage region. Furthermore, as numerical errors increase with computational load, the range of input noise that produces ideal coverage decreases with increasing FFT order. At sufficiently high FFT orders, visually beyond FFT order $25$ for the reverse transformation, ideal coverage may no longer be achievable. Although FFT is widely regarded as one of the most robust numerical algorithms, and generally insensitive to input errors, it can still fail due to numerical errors in the Library sine function. Such deterioration in calculation accuracy is not easily detectable when using conventional floating-point arithmetic.
In contrast, for Linear signals computed using the Quart sine functions, Figure 58 and 59 show the results for the forward and reverse transformations, respectively. The ideal coverage region is significantly larger, and proper coverage is also achieved in adjacent regions. The forward transformation exhibits a broader ideal coverage region than the reverse transformations. While the forward transformation shows consistent over-estimation of the value errors in the proper coverage region, the reverse transformation demonstrates both over-estimation and under-estimation of the value errors in the proper coverage region.
@@
Adding noises has another obvious benefit: Different implementations of the variance arithmetic have almost identical result when the input noise is larger than $10^{-16}$. For example, Figure 58 and 59 are almost identical using either C++, or Java, or Python.
9.13 Input Uncertainty Recovery
Figure 60: Error deviation (z-axis) versus input uncertainty (x-axis) and FFT order (y-axis) for the roundtrip transformations of Linear signals computed using the Library sine function.
Figure 61: Error deviation (z-axis) versus input uncertainty (x-axis) and FFT order (y-axis) for the roundtrip transformations of Linear signals computed using the Library sine function.
With input uncertainty unspecified, the value error waveform for the roundtrip transformation is qualitatively similar to that of the the reverse transformation, with error deviations close to $1$, such as in Figure 38, 40, 44, and 50, even when proper coverage can no longer achieve for both forward and reverse transformations in Figure 52. With input uncertainty unspecified, roundtrip transformation can always recover input uncertainty statistically.
With input uncertainty specified, roundtrip transformation can always recover input uncertainty numerically, and the quality of the recover increases with the quality of input uncertainty specification. For example, when using the Library sine function:
  • Figure 54 shows that the error deviations for the roundtrip transformation is less than $1$, suggesting that each input uncertainty is recovered numerically. These error deviations increases with FFT orders, because the error deviations without added noises increases with the FFT orders according to Figure 52.
  • As a comparison, for Sin/Cos signal, because Figure 42 shows little increase with the FFT orders of the error deviations without added noises, the improvement of the numerical recovering in Figure 61 is almost independent of FFT orders.
In both cases, error deviations decreases with the strength of the added noise.
@@ 9.14 Prec Sine Function
The Quart sine function slightly overestimates the uncertainty, while the Library sine function underestimates it significantly. This raises the question of whether a sine function with a more accurate uncertainty quantification can be developed. If the Quart sine function is constructed using the 128-bit floating-point library of gcc and then converted to 64-bit values, the difference between each 128-bit value and its corresponding 64-bit value can be used as the uncertainty for the 64-bit representation. The resulting indexed sine function is referred to as the Prec sine function.
Figure 62 compares the value errors and the uncertainties of $\sin(x)^{2}+\cos(x)^{2}-1$ between the Prec sine function and the Quart sine function, and demonstrates that the Prec sine functions has smaller uncertainties numerically than the Quart sine functions, but larger value errors statically, presumably due to rounding errors introduced by the casting. This difference indicates that casting a 128-bit floating-point value to a 64-bit floating-point value is less accurate in the 64-bit floating-point arithmetic, and the Prec sine function is expected to under-perform the Quart sine function due to under-estimation of the value errors by the uncertainties.
For Sin/Cos signals, Figure 63 is very similar Figure 38, provided that the error deviations using Prec sine functions are about $1.7$ times larger than their counterparts using the Quart sine functions.
For Linear signals, when comparing Figure 64 and 51 for the normalized value error distributions:
  • For the forward transformations, using either Prec or Quart sine function has very similar distributions.
  • For the reverse transformations, using Prec sine function results in wider distributions that using Quart sine function.
Although the ideal coverage regions for the reverse transformations remains identical in Figure 65 when compared with Figure 59, using Prec sine function results in larger error deviations in the proper coverage regions. Thus, using Prec sine function is slightly inferior than using Quart sine function. In another word, casting values from higher floating-point representation to a lower floating-point arithmetic is not necessarily superior.
9.15 Calculation Response
Figure 66: The FFT execution time (y-axis) versus the FFT order (x-axis) for the implementation using either C++, Java, or Python (legend), with each calculation consisted of $36$ roundtrip FFT transformations. The C++ executable is complied with -O3 -march=native optimization. Without such optimization, the C++ -O0 implementation is about a order of magnitude slower than the Java implementation.
Figure 66 plots the execution time for the roundtrip FFT calculations using either C++, Java, or Python. As expected, in Figure 66, the execution time increases exponentially with the FFT order $L$ as $2^{L}$ for all the three implementations when the FFT order $L>10$, validating the calculation response.
At first, it is surprising that in Figure 66 the Java implementation is 3 times faster than the C++ implementation which has been fully optimized for speed. Because FFT contains a large amount of independent and logically identical branches, the dynamic just-in-time Java compiler [37] can optimize according to data in real-time, compile the bytecodes into native machine codes, and parallelize the execution to the multiple cores and GPUs on a modern computer. Perhaps it is necessary to rethink multi-threading as a way to manually increase the time performance, because the dynamic just-in-time optimization of Java is superior to the static optimization of C++ for repeated and heavy computations with preload data such as FFT.
@@ 9.16 Summary
The Library sine functions implemented using conventional floating-point arithmetic have been shown to contain numerical errors equivalent up to $10^{-3}$ of the input precision in the worst cases of FFT transformations. These $\sin(x)$ numerical errors increase periodically with $x$, potentially resonating with periodic input signals. Computational result obtained with floating-point arithmetic are highly sensitive to the input data, the scale of computation, and the numerical errors in library functions. This implies that a small-scale test cannot reliably qualify or predict the behavior of large-scale calculations, but needs rigorous analysis for the ideal or proper coverage regions in each parameter space. The impact of numerical errors within mathematical libraries has not received sufficient attention. Moreover, such errors in floating-point arithmetic are inherently difficult to detect and quantify.
@@ 11.3 Improvements Needed
Bound momentum $\zeta(2n)$ needs to be extended to all probability distributions. Its asymptotic behavior $2n\rightarrow+\infty:\zeta(2n)\rightarrow\kappa^{2n}/(2n)$ needs to be generalized. Perhaps the asymptotic behavior implies a generic new requirements on the probability distributions.
The choice of ideal bounding range needs further study. For example, the choice of $\kappa_{s}=5$ for Normal distribution is based on the $5-\sigma$ rule in practice only. Such choice certainly can not be applied to other distributions.
@@
The dependence of the convergence of $\log(x\pm\delta x)$ and $(x\pm\delta x)^{c}$ on the chose of ideal bounding range $\kappa$ needs further discussion. Determining the uncertainty upper bounds for each analytic expression through analytic methods is an important next step.
@@
As an enhancement to dependency tracing, source tracing identifies each input’s contribution to the overall result uncertainty. This capability enables engineers to pinpoint the primary sources of measurement inaccuracy and in turn guide targeted improvements in data acquisition and processing strategies. For example, Formula (2.50) can guide how to improve the ideal ratio of $x\pm y$.
A key open question is whether variance arithmetic can be adapted to achieve ideal coverage for floating-point rounding errors. Except for integer multiplication, variance arithmetic does not track rounding errors during computation, because tracking rounding errors for every operation is too expensive to achieve in software. The sensitivity to the library functions increases the difficulty to solve these types of problems. Developing variance arithmetic with ideal coverage for such errors would be quite valuable because many theoretical calculations lack explicit input uncertainties.
@@ 12.1 Acknowledgments
12.2 Data Availability Statement
The data set used in this study are all generated in the open-source project at https://github.com/Chengpu0707/VarianceArithmetic. The execution assistance and explanation of the above code are available from the author upon reasonable request.
12.3 Competing Interests
The author has no competing interests to declare that are relevant to the content of this article.
12.4 Founding
No funding was received from any organization or agency in support of this research.